OpenStudy (anonymous):

HELP ME PLEASE FAN/MEDAL Write the first 5 terms of a geometric sequence. Then find the 10th and 20th terms by using the formula.

OpenStudy (imstuck):

you could use something simple and basic like {1,3,9,27,81,...}. The ratio is 3. so the formula would be \[a _{n}=1\times (3)^{n-1}\]

OpenStudy (imstuck):

The 10th term, then, would be\[a _{10}=1\times (3)^{10-1}\]which is 19683, and for the 20th term, you would have\[a _{20}=1\times (3)^{20-1}\]which would be 1162261467.

OpenStudy (anonymous):

AWESOME THANK YOU!!!!!!!!!!